OpenStudy (melstutes):

What is the equation of the line parallel to the line 5x-2x = 7 through the ordered pair (6, 0)?

OpenStudy (shadowreaper810):

The equation is 5x-2x=7 This is equal to 3x=7 => x=7/3 this equation is parallel to the y axis as it is in the form x=a now the other equation that is parallel to it is also parallel to y axis and will be in the form x=a We know that the x cordinate is supposed to be 6 as the mentioned in question that the line should pass through(6,0) so x=6 and this is the you required equation. this is the answer.

OpenStudy (anonymous):

to have a line you need 2 variables, f(x,y)=5x-2y=7 So a parallel line to that must have same slope, diferent point \[f_{//} (x,y) = 5x-2y = something\ called:\ contour\ line\ "C"\] So \[f(6,0) = 5·6 -2·0 =C \] Then C=30 SOOOO your new line is f(x,y)=5x-2y=30 ta da! Note. the slope is the number next to each variable. 5 for X, and -2 for Y Really is \[5=d f(x,y)/dx \ \ ,\ -2=d f(x,y)/dy\] i'm trying this page for first (maybe unique) time, i put a random question, don't expect a response :P